# Forecasting high-impact research topics via machine learning on evolving knowledge graphs

### v3

@@ Results
To build our concept list, we started with 2,444,442 papers from four publicly available preprint servers: arXiv, bioRxiv, medRxiv, and chemRxiv. We use papers from preprint servers for two reasons: (1) It contains papers that are not published yet in journals, thus our dataset also contains state-of-the-art concepts; (2) they associate papers to research categories, which can be used to focus on specific scientific domains (as we do, with the field of quantum physics and optics). The data cutoff is February 2023. From these, we extracted titles and abstracts of the papers. To single out concept candidates from this extensive collection, we applied the Rapid Automatic Keyword Extraction (RAKE) algorithm based on statistical text analysis to automatically detect important keywords [41]. Concepts with two words, like phase transition, were retained if they appear in at least 9 papers, while longer concepts, such as single molecule localization microscopy, needed to appear in at least 6 papers. In this way, we can increase the fraction of high-quality concepts. We further developed a suite of natural language processing tools to refine the concepts, followed by manual inspection to remove any incorrectly identified ones. Finally, we got a list which contains over 368,000 concepts. We focus here on concepts specific to the sub-field of optics and quantum physics (representing roughly 10% of the entire concepts), but our method can immediately be translated to any other domain. This refined domain-specific concept list serves as the vertices of our knowledge graph.
@@
Figure 5: Network features vs. predicted impact. A randomly selected set of 100,000 unconnected pairs until January 2023 is used. The color represents the neural network’s prediction of each concept pair’s impact. (a): The y-axis shows cosine similarity, indicating the semantic similarity between concepts; lower values represent concept pairs that are semantically distinct. The x-axis is the average vertex degree of the two concepts in the knowledge graph, reflecting their overall prominence. Concepts with low similarity and low degree yet predicted to have high impact could be surprising and offer interesting suggestions. (b): The x-axis represents the average number of new neighbors each concept gained over the last three years. Concept pairs with low similarity and few new neighbors but high impact predictions might highlight potentially overlooked but intriguing ideas. (c): The x-axis denotes citation density (average citations per paper mentioning the concepts). Pairs with low similarity and citation density but high predicted impact could again indicate overlooked potential ideas. (d): Citation counts for concept 1 (x-axis) and concept 2 (y-axis) over last three years are plotted on a logarithmic scale. We can easily identify concept pairs predicted to have high impact in the future, even though they have individually received few citations in the past.
@@
The final constructed knowledge graph has 37,960 vertices with more than 26 million edges (built from 190 million concept pairs, containing multi-edges when multiple papers create the same edge) from the OpenAlex dataset, with a data cutoff at April 2023. In Fig. 2, we show the fastest growing (in terms of citation) concepts and concept pairs since 2012, where we can recognize many highly influential topics in quantum physics and optics research.
Forecasting impact of newly created concept connections – With an evolving knowledge graph from the past, we can formulate the prediction of impact for new concept pairs as a supervised learning task, as illustrated in Fig. 3. For a vertex pair that has not had any connection in the year 2016, we predict whether three years later this pair accumulated more than a certain number of citations. Using the historical knowledge graph, we possess an ideal supervision signal for our binary classification task. During the training phase, we selected pairs of vertices that were not connected and calculated 141 features for each pair. These features include 41 network features, divided into 20 node features (such as the number of neighbors and PageRank [43] over the past three years) and 21 edge features (including cosine, geometric, and Simpson similarities [44]). Additionally, we incorporated 100 impact features: 58 of these are node citation features, covering total citations and yearly citations within the last three years. The other 42 features are about vertex pairs and include measures such as the citation ratio between them. Detailed feature description are available in GitHub: Impact4Cast. The network features are inspired by the winner of the Science4Cast competition [45, 11], and the citation features are developed empirically and could potentially be improved by careful feature importance analysis. Our neural network is a fully connected feed-forward network with four hidden layers of 600 neurons each. The exploration of more advanced architectures might improve the prediction qualities further. The neural network has to predict whether the unconnected vertex pair in 2019 will have at least $IR$ citations ($IR$ stands for the impact range).
@@ Discussion
We show how to forecast the impact of future research topics. Although we view this as a significant step towards developing truly useful AI-driven assistants, achieving this goal requires numerous further advancements. Firstly, developing methods to extract more complex information from each paper will be crucial, for instance by employing hyper-graph structures that carry more information from each paper [55, 56], which has already been demonstrated to lead to exciting results in other domains [57, 2, 58]. This might also allow for the forecast of new concepts [59, 60] and their impact. Incorporating the recent dataset [61, 62] into our research could also allow us to explore more complex data structures than those used in our paper. Secondly, it will be interesting to approximate impact with metrics that go beyond citations – which is a crucial topic in computational sociology and the study of the science of science [1, 2]. Additionally, introducing metrics of surprise, as discussed in [63, 64], could serve as a complementary metric to citation prediction for ranking suggestions. Finally, while the suggestion of impactful new ideas might be a key component of future AI assistants, it will be crucial to study its relation to the scientific interest of working researchers [65].
Materials and Methods
Datasets for knowledge graph construction – To compile a list of scientific concepts in natural science, we used metadata from four major preprint servers: arXiv, bioRxiv, medRxiv, and chemRxiv. The arXiv dataset can be directly downloaded from Kaggle, while metadata from bioRxiv, medRxiv, and chemRxiv are accessible through their APIs. The full methodology and codes are available on the GitHub: Impact4Cast. Our comprehensive dataset encompasses approximately 2.44 million papers, including 78,084 from arXiv’s physics.optics and quant-ph categories, which were specifically utilized for identifying domain concepts.
@@
Details on concept and edge generation – From the preprint dataset of $\sim$2.44 million papers, we analyzed each article’s title and abstract using the RAKE algorithm, enhanced with additional stopwords, to extract potential concept candidates. These candidates were stored for subsequent analysis. We filtered out concepts to retain only those with two words that appeared in nine or more articles, and those with three or more words that appeared in six or more articles. This step significantly reduced the noise from the RAKE-generated concepts, yielding a refined list of 726,439 relevant concepts. To further enhance the quality of the identified concepts, we developed a suite of automatic tools designed to identify and eliminate common, domain-independent errors often associated with RAKE. In addition, we conduct a manual review to identify and eliminate any inaccuracies in the concepts. The entire process, which included eliminating non-conceptual phrases, verbs, ordinal numbers, conjunctions, and adverbials, resulted in a full list of 368,825 concepts.
@@
Training – Our neural network consists of six fully connected layers, which include four hidden layers with 600 neurons each. The network inputs are 141 features for each unconnected concept pair $(v_{i},v_{j})$, denoted as $p_{i,j}=(p^{1}_{i,j},p^{2}_{i,j},...,p^{141}_{i,j})$, where each $p_{i,j}\in\mathbb{R}$. For instance, $p^{1}_{i,j}$ and $p^{2}_{i,j}$ represent the vertex degree of concepts $v_{i}$ and $v_{j}$ for the current year, $y$. Detailed feature description and feature generation code are available in GitHub: Impact4Cast.
@@
Individual feature’s predictive ability – In Fig. 4 (a), we observe an AUC score of 0.948 for the 2019 evaluation dataset with the neural network that uses all 141 features, trained on 2016 dataset and impact range $IR=100$. To explore the predictive ability of individual features, we trained separate neural networks on each feature using the same 2016 dataset, and then applied the 2019 evaluation dataset to these models. This resulted in 141 individual predictions, each from a network trained on a single feature. The features were ranked by their impact predictions, shown in the Fig. 12.
Acknowledgements
Data and Code availability: Data is accessible on Zenodo at https://doi.org/10.5281/zenodo.10692137 [66]. Benchmark data used in our work is also available on Zenodo at https://doi.org/10.5281/zenodo.14527306 [67]. Codes for this work are available on GitHub at https://github.com/artificial-scientist-lab/Impact4Cast.

### v4

@@ Results
To build our concept list, we started with 2,444,442 papers from four publicly available preprint servers: arXiv, bioRxiv, medRxiv, and chemRxiv. We use papers from preprint servers for two reasons: (1) It contains papers that are not published yet in journals, thus our dataset also contains state-of-the-art concepts; (2) they associate papers to research categories, which can be used to focus on specific scientific domains (as we do, with the field of quantum physics and optics). The data cutoff is February 2023. From these, we extracted titles and abstracts of the papers. To single out concept candidates from this extensive collection, we applied the Rapid Automatic Keyword Extraction (RAKE) algorithm based on statistical text analysis to automatically detect important keywords [41] (see details in the Appendix B). Each of these candidates are ranked Concepts with two words, like phase transition, were retained if they appear in at least 9 papers, while longer concepts, such as single molecule localization microscopy, needed to appear in at least 6 papers. In this way, we can increase the fraction of high-quality concepts. We further developed a suite of natural language processing tools to refine the concepts, followed by manual inspection to remove any incorrectly identified ones. Finally, we got a list which contains over 368,000 concepts. We focus here on concepts specific to the sub-field of optics and quantum physics (representing roughly 10% of the entire concepts), but our method can immediately be translated to any other domain. This refined domain-specific concept list serves as the vertices of our knowledge graph.
@@
Figure 5: Network features vs. predicted impact. A randomly selected set of 100,000 unconnected pairs until January 2023 is used. The color represents the neural network’s prediction of each concept pair’s impact. (a): The y-axis shows cosine similarity, indicating the semantic similarity between concepts; lower values represent concept pairs that are semantically distinct. For each concept pair $(u,v)$, cosine similarity is calculated as the number of shared neighbors divided by the square root of the product of the number of neighbors of $u$ and $v$. The x-axis is the average vertex degree of the two concepts in the knowledge graph, reflecting their overall prominence. Concepts with low similarity and low degree yet predicted to have high impact could be surprising and offer interesting suggestions. (b): The x-axis represents the average number of new neighbors each concept gained over the last three years. Concept pairs with low similarity and few new neighbors but high impact predictions might highlight potentially overlooked but intriguing ideas. (c): The x-axis denotes citation density (average citations per paper mentioning the concepts). Pairs with low similarity and citation density but high predicted impact could again indicate overlooked potential ideas. (d): Citation counts for concept 1 (x-axis) and concept 2 (y-axis) over last three years are plotted on a logarithmic scale. We can easily identify concept pairs predicted to have high impact in the future, even though they have individually received few citations in the past.
@@
The final constructed knowledge graph has 37,960 vertices with more than 26 million edges (built from 190 million concept pairs, containing multi-edges when multiple papers create the same edge) from the OpenAlex dataset, with a data cutoff at April 2023.
In Fig. 2, we show the fastest growing (in terms of citation) concepts and concept pairs since 2012, where we can recognize many highly influential topics in quantum physics and optics research.
Forecasting impact of newly created concept connections – With an evolving knowledge graph from the past, we can formulate the prediction of impact for new concept pairs as a supervised learning task, as illustrated in Fig. 3. For a vertex pair that has not had any connection in the year 2016, we predict whether three years later this pair accumulated more than a certain number of citations. Using the historical knowledge graph, we possess an ideal supervision signal for our binary classification task. During the training phase, we selected pairs of vertices that were not connected and calculated 141 features for each pair. These features include 41 network features, divided into 20 node features (such as the number of neighbors and PageRank [43] over the past three years) and 21 edge features (including cosine, geometric, and Simpson similarities [44]). Additionally, we incorporated 100 impact features: 58 of these are node citation features, covering total citations and yearly citations within the last three years. The other 42 features are about vertex pairs and include measures such as the citation ratio between them. Detailed feature description are available in GitHub: Impact4Cast and Tables. 1 and  2 in Appendix D. The network features are inspired by the winner of the Science4Cast competition [45, 11], and the citation features are developed empirically and could potentially be improved by careful feature importance analysis. Our neural network is a fully connected feed-forward network with four hidden layers of 600 neurons each. The exploration of more advanced architectures might improve the prediction qualities further. The neural network has to predict whether the unconnected vertex pair in 2019 will have at least $IR$ citations ($IR$ stands for the impact range).
@@ Discussion
We show how to forecast the impact of future research topics. Although we view this as a significant step towards developing truly useful AI-driven assistants, achieving this goal requires numerous further advancements. Firstly, developing methods to extract more complex information from each paper will be crucial, for instance by employing hyper-graph structures that carry more information from each paper [55, 56], which has already been demonstrated to lead to exciting results in other domains [57, 2, 58]. The forecast itself could benefit from more genuine dynamical features that go beyond network snapshots from different years [59], or the application of dynamic word embedding [60]. This might also allow for the forecast of new concepts [61, 62] and their impact. Incorporating the recent dataset [63, 64] into our research could also allow us to explore more complex data structures than those used in our paper. Secondly, it will be interesting to approximate impact with metrics that go beyond citations – which is a crucial topic in computational sociology and the study of the science of science [1, 2]. Additionally, introducing metrics of surprise, as discussed in [65, 66], could serve as a complementary metric to citation prediction for ranking suggestions. Finally, while the suggestion of impactful new ideas might be a key component of future AI assistants, it will be crucial to study its relation to the scientific interest of working researchers [67].
Acknowledgements
Data availability statement Data is accessible on Zenodo at https://doi.org/10.5281/zenodo.10692137 [68]. Benchmark data used in our work is also available on Zenodo at https://doi.org/10.5281/zenodo.14527306 [69]. Codes for this work are available on GitHub at https://github.com/artificial-scientist-lab/Impact4Cast.
Appendix A Datasets for the knowledge graph
To compile a list of scientific concepts in natural science, we used metadata from four major preprint servers: arXiv, bioRxiv, medRxiv, and chemRxiv. The arXiv dataset can be directly downloaded from Kaggle, while metadata from bioRxiv, medRxiv, and chemRxiv are accessible through their APIs. The full methodology and codes are available on the GitHub: Impact4Cast. Our comprehensive dataset encompasses approximately 2.44 million papers, including 78,084 from arXiv’s physics.optics and quant-ph categories, which were specifically utilized for identifying domain concepts.
@@ Appendix A Datasets for the knowledge graph
Appendix B Details on concept and edge generation
From the preprint dataset of $\sim$2.44 million papers, we analyzed each article’s title and abstract using the RAKE algorithm.
RAKE works by proposing candidate key phrases from each sentence by splitting the sentence at punctuation marks and stop words. Let’s take the sentence: Recurrent neural networks have significantly improved the accuracy of image recognition in large-scale scientific collaboration networks. RAKE splits this into the candidate phrases Recurrent neural networks, improved, accuracy, image recognition, and large-scale scientific collaboration networks. Each individual word within these candidate phrases receives a score based on two metrics: its frequency (how often it occurs in candidate phrases; here only networks occurs twice, all others once) and its degree (the number of co-occurrences with other words in candidate phrases). For example, the word networks has a higher degree because it appears in multiple longer phrases, thus increasing its importance. The final keyword phrases are ranked based on the summed scores of their individual words, naturally favoring longer, meaningful phrases such as large-scale scientific collaboration networks or recurrent neural networks.
RAKE’s candidates are then used for subsequent analysis. We filtered out concepts to retain only those with two words that appeared in nine or more articles, and those with three or more words that appeared in six or more articles. This step significantly reduced the noise from the RAKE-generated concepts, yielding a refined list of 726,439 relevant concepts. To further enhance the quality of the identified concepts, we developed a suite of automatic tools designed to identify and eliminate common, domain-independent errors often associated with RAKE. In addition, we conduct a manual review to identify and eliminate any inaccuracies in the concepts. The entire process, which included eliminating non-conceptual phrases, verbs, ordinal numbers, conjunctions, and adverbials, resulted in a full list of 368,825 concepts.
We note that our approach might result in similar or synonymous concepts appearing as separate key phrases due to slight variations or phrasing differences. Although multiple entries in the extracted concept list may represent essentially the same idea, synonymity can effectively be identified by leveraging structural information within the knowledge graph. Specifically, network cosine similarity between nodes (concepts) reveals the degree of similarity in their network neighborhoods, indicating semantic closeness. Practically, this can be utilized to enhance impact forecasting: when recommending new high-impact concept pairs, applying a cosine similarity threshold ensures that suggested pairs have sufficiently distinct meanings, thus avoiding redundancy from synonymous phrases.
@@ Appendix B Details on concept and edge generation
Appendix C Details on training
Our neural network consists of six fully connected layers, which include four hidden layers with 600 neurons each. The network inputs are 141 features for each unconnected concept pair $(v_{i},v_{j})$, denoted as $p_{i,j}=(p^{1}_{i,j},p^{2}_{i,j},...,p^{141}_{i,j})$, where each $p_{i,j}\in\mathbb{R}$. For instance, $p^{1}_{i,j}$ and $p^{2}_{i,j}$ represent the vertex degree of concepts $v_{i}$ and $v_{j}$ for the current year, $y$. Detailed feature description and feature generation code are available in GitHub: Impact4Cast.
@@ Appendix D Individual feature’s predictive ability
In Fig. 4 (a), we observe an AUC score of 0.948 for the 2019 evaluation dataset with the neural network that uses all 141 features, trained on 2016 dataset and impact range $IR=100$. To explore the predictive ability of individual features, we trained separate neural networks on each feature using the same 2016 dataset, and then applied the 2019 evaluation dataset to these models. This resulted in 141 individual predictions, each from a network trained on a single feature. The features were ranked by their impact predictions, shown in the Fig. 12. Details of the features are shown in Tables 1 and  1, and the corresponding documentation and source code are available at GitHub: Impact4Cast.
Table 1: Detailed description of all customized features for an unconnected concept pair $(u,v)$ for the year $y$
Type Feature Index Description
node feature 0-5 Number of neighbors for each node ($u$ or $v$) until the year $y$, $y{-}1$, $y{-}2$ denoted as $N_{u,y}$, $N_{v,y}$, $N_{u,y-1}$, $N_{v,y-1}$, $N_{u,y-2}$, and $N_{v,y-2}$, ordered as indices 0–5.
6-7 Number of new neighbors for each node ($u$ or $v$) between year $y{-}1$ and $y$ i.e., $N_{u,y}{-}N_{u,y-1}$ and $N_{v,y}{-}N_{v,y-1}$.
8-9 Number of new neighbors for each node ($u$ or $v$) between year $y{-}2$ and $y$ i.e., $N_{u,y}{-}N_{u,y-2}$ and $N_{v,y}{-}N_{v,y-2}$.
10-11 Rank of the number of new neighbors for each node ($u$ or $v$) between year $y{-}1$ and $y$ i.e., rank($N_{u,y}{-}N_{u,y-1}$) and rank($N_{v,y}{-}N_{v,y-1}$).
12-13 Rank of the number of new neighbors for each node ($u$ or $v$) between year $y{-}2$ and $y$ i.e., rank($N_{u,y}{-}N_{u,y-1}$) and rank($N_{v,y}{-}N_{v,y-2}$).
14-19 PageRank scores of each node ($u$ or $v$) until the year $y$, $y{-}1$, $y{-}2$ denoted and ordered as $\mathrm{PR}_{u,y}$, $\mathrm{PR}_{v,y}$, $\mathrm{PR}_{u,y-1}$, $\mathrm{PR}_{v,y-1}$, $\mathrm{PR}_{u,y-2}$ and $\mathrm{PR}_{v,y-2}$.
node citation feature 20-25 Yearly citation for each node ($u$ or $v$) in year $y$, $y{-}1$, $y{-}2$ denoted and ordered as $\mathrm{C}_{u,y}$, $\mathrm{C}_{v,y}$, $\mathrm{C}_{u,y-1}$, $\mathrm{C}_{v,y-1}$, $\mathrm{C}_{u,y-2}$ and $\mathrm{C}_{v,y-2}$.
26-31 Total citation for each node ($u$ or $v$) since the first publication to the year $y$, $y{-}1$, and $y{-}2$ denoted and ordered as $\mathrm{Ct}_{u,y}$, $\mathrm{Ct}_{v,y}$, $\mathrm{Ct}_{u,y-1}$, $\mathrm{Ct}_{v,y-1}$, $\mathrm{Ct}_{u,y-2}$ and $\mathrm{Ct}_{v,y-2}$.
32-37 Total citations for each node ($u$ or $v$) in the last 3 years ending in the year $y$, $y{-}1$, and $y{-}2$ denoted and ordered as $\mathrm{Ct}^{\Delta 3}_{u,y}$, $\mathrm{Ct}^{\Delta 3}_{v,y}$, $\mathrm{Ct}^{\Delta 3}_{u,y{-}1}$, $\mathrm{Ct}^{\Delta 3}_{v,y{-}1}$, $\mathrm{Ct}^{\Delta 3}_{u,y{-}2}$, and $\mathrm{Ct}^{\Delta 3}_{v,y{-}2}$.
38-43 Number of papers mentioning node $u$ from the first publication to the year $y$, $y{-}1$, and $y{-}2$, similar for node $v$; denoted and ordered as $\mathrm{Pn}_{u,y}$, $\mathrm{Pn}_{v,y}$, $\mathrm{Pn}_{u,y-1}$, $\mathrm{Pn}_{v,y-1}$, $\mathrm{Pn}_{u,y-2}$, and $\mathrm{Pn}_{v,y-2}$
44-49 Average yearly citations for each node ($u$ or $v$) in the year $y$, $y{-}1$, $y{-}2$ denoted and ordered as $\mathrm{Cm}_{u,y}$, $\mathrm{Cm}_{v,y}$, $\mathrm{Cm}_{u,y-1}$, $\mathrm{Cm}_{v,y-1}$, $\mathrm{Cm}_{u,y-2}$ and $\mathrm{Cm}_{v,y-2}$ e.g., $\mathrm{Cm}_{u,y}=\mathrm{C}_{u,y}/\mathrm{Pn}_{u,y}$
50-55 Average total citations for each node ($u$ or $v$) since the first publications to the years $y$, $y{-}1$, $y{-}2$; denoted and ordered as $\mathrm{Ctm}_{u,y}$, $\mathrm{Ctm}_{v,y}$, $\mathrm{Ctm}_{u,y-1}$, $\mathrm{Ctm}_{v,y-1}$, $\mathrm{Ctm}_{u,y-2}$ and $\mathrm{Ctm}_{v,y-2}$; e.g., $\mathrm{Ctm}_{u,y}=\mathrm{Ct}_{u,y}/\mathrm{Pn}_{u,y}$
56-61 Average total citations for each node ($u$ or $v$) in the last 3 years ending in the year $y$, $y{-}1$, and $y{-}2$; denoted and ordered as $\mathrm{Ctm}^{\Delta 3}_{u,y}$, $\mathrm{Ctm}^{\Delta 3}_{v,y}$, $\mathrm{Ctm}^{\Delta 3}_{u,y-1}$, $\mathrm{Ctm}^{\Delta 3}_{v,y-1}$, $\mathrm{Ctm}^{\Delta 3}_{u,y-2}$ and $\mathrm{Ctm}^{\Delta 3}_{v,y-2}$; e.g., $\mathrm{Ctm}^{\Delta 3}_{u,y}=\mathrm{Ct}^{\Delta 3}_{u,y}/\mathrm{Pn}_{u,y}$
62-63 New citations for each node ($u$ or $v$) between years $y{-}1$ and $y$ i.e., $\mathrm{Ct}_{u,y}{-}\mathrm{Ct}_{u,y-1}$ and $\mathrm{Ct}_{v,y}{-}\mathrm{Ct}_{v,y-1}$.
64-65 New citations for each node ($u$ or $v$) between years $y{-}2$ and $y$ i.e., $\mathrm{Ct}_{u,y}{-}\mathrm{Ct}_{u,y-2}$ and $\mathrm{Ct}_{v,y}{-}\mathrm{Ct}_{v,y-2}$.
66-67 Rank of the new citations for each node ($u$ or $v$) between years $y{-}1$ and $y$ i.e., rank($\mathrm{C}_{u,y}{-}\mathrm{C}_{u,y-1}$) and rank($\mathrm{C}_{v,y}{-}\mathrm{C}_{v,y-1}$).
68-69 Rank of the new citations for each node ($u$ or $v$) between years $y{-}2$ and $y$ i.e., rank($\mathrm{C}_{u,y}{-}\mathrm{C}_{u,y-2}$) and rank($\mathrm{C}_{v,y}{-}\mathrm{C}_{v,y-2}$).
70-71 Number of papers mentioning nodes $u$ between years $y{-}1$ and $y$, similar for node $v$ i.e., $\mathrm{PR}_{u,y}-\mathrm{PR}_{u,y-1}$ and $\mathrm{PR}_{v,y}-\mathrm{PR}_{v,y-1}$
72-73 Number of papers mentioning nodes $u$ between years $y{-}2$ and $y$, similar for node $v$ i.e., $\mathrm{PR}_{u,y}-\mathrm{PR}_{u,y-2}$ and $\mathrm{PR}_{v,y}-\mathrm{PR}_{v,y-2}$
74-75 Rank of the number of papers mentioning nodes $u$ between years $y{-}1$ and $y$, similar for node $v$; i.e., rank($\mathrm{PR}_{u,y}-\mathrm{PR}_{u,y-1}$) and rank($\mathrm{PR}_{v,y}-\mathrm{PR}_{v,y-1}$)
76-77 Number of papers mentioning nodes $u$ between years $y{-}2$ and $y$, similar for node $v$ i.e., rank($\mathrm{PR}_{u,y}-\mathrm{PR}_{u,y-2}$) and rank($\mathrm{PR}_{v,y}-\mathrm{PR}_{v,y-2}$)
Table 2: continued from previous page for Table. 1
Type Feature Index Description
pair feature 78-80 Number of shared neighbors between nodes $u$ and $v$ until the year $y$, $y{-}1$, $y{-}2$, denoted and ordered as $\mathrm{Ns}_{y}$, $\mathrm{Ns}_{y-1}$ and $\mathrm{Ns}_{y-2}$; e.g., $\mathrm{Ns}_{y}=\mathrm{N}_{u,y}\cap\mathrm{N}_{v,y}$
81-83 Geometric similarity coefficient for the pair $(u,v)$ for the year $y$, $y{-}1$, and $y{-}2$ denoted and ordered as $\mathrm{Geo}_{y}$, $\mathrm{Geo}_{y-1}$, and $\mathrm{Geo}_{y-2}$; e.g., $\mathrm{Geo}_{y}=\mathrm{Ns}_{y}^{2}/(\mathrm{N}_{u,y}\times\mathrm{N}_{v,y})$.
84-86 Cosine similarity coefficient for the pair $(u,v)$ for the year $y$, $y{-}1$, $y{-}2$ denoted and ordered as $\mathrm{Cos}_{y}$, $\mathrm{Cos}_{y-1}$, and $\mathrm{Cos}_{y-2}$; e.g., $\mathrm{Cos}_{y}=\sqrt{\mathrm{Geo}_{y}}$.
87-89 Simpson coefficient for the pair $(u,v)$ for the year $y$, $y{-}1$, $y{-}2$ denoted and ordered as $\mathrm{Sim}_{y}$, $\mathrm{Sim}_{y-1}$, and $\mathrm{Sim}_{y-2}$; e.g., $\mathrm{Sim}_{y}=\mathrm{Ns}_{y}/\min(\mathrm{N}_{u,y},\mathrm{N}_{v,y})$.
90-92 Preferential attachment coefficient for the pair $(u,v)$ for the year $y$, $y{-}1$, $y{-}2$ denoted and ordered as $\mathrm{Pre}_{y}$, $\mathrm{Pre}_{y-1}$, and $\mathrm{Pre}_{y-2}$; e.g., $\mathrm{Pre}_{y}=\mathrm{N}_{u,y}\times\mathrm{N}_{v,y}$.
93-95 Sørensen–Dice coefficient for the pair $(u,v)$ for the year $y$, $y{-}1$, $y{-}2$ denoted and ordered as $\mathrm{Sor}_{y}$, $\mathrm{Sor}_{y-1}$, and $\mathrm{Sor}_{y-2}$; e.g., $\mathrm{Sor}_{y}=2\mathrm{Ns}_{y}/(\mathrm{N}_{u,y}+\mathrm{N}_{v,y}$).
96-98 Jaccard coefficient for the pair $(u,v)$ for the year $y$, $y{-}1$, $y{-}2$ denoted and ordered as $\mathrm{Jac}_{y}$, $\mathrm{Jac}_{y-1}$, and $\mathrm{Jac}_{y-2}$; e.g., $\mathrm{Jac}_{y}=\mathrm{Ns}_{y}/(\mathrm{N}_{u,y}+\mathrm{N}_{v,y}-\mathrm{Ns}_{y})$.
[…]
105-107 Sum of average citations received by nodes $u$ and $v$ in the year $y$, $y{-}1$, and $y{-}2$ denoted and ordered as $\mathrm{s}_{y}$, $\mathrm{s}_{y-1}$, and $\mathrm{s}_{y-2}$; e.g., $\mathrm{s}_{y}=\mathrm{Cm}_{u,y}+\mathrm{Cm}_{v,y}$.
108-110 Sum of average total citations received by nodes $u$ and $v$ from the first publication to the year $y$, $y{-}1$, and $y{-}2$; denoted and ordered as $\mathrm{st}_{y}$, $\mathrm{st}_{y-1}$, and $\mathrm{st}_{y-2}$; e.g., $\mathrm{st}_{y}=\mathrm{Ctm}_{u,y}+\mathrm{Ctm}_{v,y}$.
111-113 Sum of the total citations received by nodes $u$ and $v$ in the last 3 years ending in the year $y$, $y{-}1$, and $y{-}2$; denoted and ordered as $\mathrm{st}^{\Delta 3}_{y}$, $\mathrm{st}^{\Delta 3}_{y-1}$, and $\mathrm{st}^{\Delta 3}_{y-2}$; e.g., $\mathrm{st}^{\Delta 3}_{y}=\mathrm{Ct}^{\Delta 3}_{u,y}+\mathrm{Ct}^{\Delta 3}_{v,y}$.
114-116 Sum of average total citations received by nodes $u$ and $v$ in the last 3 years ending in the year $y$, $y{-}1$, and $y{-}2$; denoted and ordered as $\mathrm{stm}^{\Delta 3}_{y}$, $\mathrm{stm}^{\Delta 3}_{y-1}$, and $\mathrm{stm}^{\Delta 3}_{y-2}$ e.g., $\mathrm{stm}^{\Delta 3}_{y}=\mathrm{Ctm}^{\Delta 3}_{u,y}+\mathrm{Ctm}^{\Delta 3}_{v,y}$.
117-119 Minimum number of citations received by either node $u$ or $v$ in years $y$, $y{-}1$, and $y{-}2$; denoted and ordered as $\mathrm{minC}_{y}$, $\mathrm{minC}_{y-1}$, and $\mathrm{minC}_{y-2}$; e.g., $\mathrm{minC}_{y}=\min(\mathrm{C}_{u,y},\mathrm{C}_{v,y})$.
120-122 Maximum number of citations received by either node $u$ or $v$ in years $y$, $y{-}1$, and $y{-}2$; denoted and ordered as $\mathrm{maxC}_{y}$, $\mathrm{maxC}_{y-1}$, and $\mathrm{maxC}_{y-2}$; e.g., $\mathrm{maxC}_{y}=\max(\mathrm{C}_{u,y},\mathrm{C}_{v,y})$.
123-125 Minimum number of total citations received by nodes $u$ and $v$ from the first publication to the year $y$, $y{-}1$, and $y{-}2$; denoted and ordered as $\mathrm{minCt}_{y}$, $\mathrm{minCt}_{y-1}$, and $\mathrm{minCt}_{y-2}$; e.g., $\mathrm{minCt}_{y}=\min(\mathrm{Ct}_{u,y},\mathrm{Ct}_{v,y})$.
126-128 Maximum number of total citations received by nodes $u$ and $v$ from the first publication to the year $y$, $y{-}1$, and $y{-}2$; denoted and ordered as $\mathrm{maxCt}_{y}$, $\mathrm{maxCt}_{y-1}$, and $\mathrm{maxCt}_{y-2}$; e.g., $\mathrm{maxCt}_{y}=\max(\mathrm{Ct}_{u,y},\mathrm{Ct}_{v,y})$.
129-131 Minimum number of total citations received by nodes $u$ and $v$ in the last 3 years ending in the year $y$, $y{-}1$, and $y{-}2$; denoted and ordered as $\mathrm{minCt}^{\Delta 3}_{y}$, $\mathrm{minCt}^{\Delta 3}_{y-1}$, and $\mathrm{minCt}^{\Delta 3}_{y-2}$; e.g., $\mathrm{minCt}^{\Delta 3}_{y}=\min(\mathrm{Ct}^{\Delta 3}_{u,y},\mathrm{Ct}^{\Delta 3}_{v,y})$.
132-134 Maximum number of total citations received by nodes $u$ and $v$ in the last 3 years ending in the year $y$, $y{-}1$, and $y{-}2$; denoted and ordered as $\mathrm{maxCt}^{\Delta 3}_{y}$, $\mathrm{maxCt}^{\Delta 3}_{y-1}$, and $\mathrm{maxCt}^{\Delta 3}_{y-2}$; e.g., $\mathrm{maxCt}^{\Delta 3}_{y}=\max(\mathrm{Ct}^{\Delta 3}_{u,y},\mathrm{Ct}^{\Delta 3}_{v,y})$.
135-137 Minimum number of papers mentioning the node $u$ or node $v$ from the first publication to the year $y$, $y{-}1$, and $y{-}2$; denoted and ordered as $\mathrm{minPn}_{y}$, $\mathrm{minPn}_{y-1}$ and $\mathrm{minPn}_{y-2}$; e.g., $\mathrm{minPn}_{y}=\min(\mathrm{Pn}_{u,y},\mathrm{Pn}_{v,y})$.
138-140 Maximum number of papers mentioning the node $u$ or node $v$ from the first publication to the year $y$, $y{-}1$, and $y{-}2$; denoted and ordered as $\mathrm{maxPn}_{y}$, $\mathrm{maxPn}_{y-1}$ and $\mathrm{maxPn}_{y-2}$; e.g., $\mathrm{maxPn}_{y}=\max(\mathrm{Pn}_{u,y},\mathrm{Pn}_{v,y})$.